1.   In an experiment, 112cm3 of carbon (iv) oxide gas were collected at s.t.p. How many molecules of the gas are present? (molar volume of the gas at stp=22.4dm, L=6.02 x 1023mol-1)



1
Expert's answer
2022-02-23T01:28:05-0500

Convert "112cm^3" of "CO_2" to "dm^3=112\times 0.001=0.112dm^3"

Molar volume of the gas at stp= "22.4dm^3=6.02\times 10^{23}mol^{-1}"

Molecules of "0.112dm^3" of carbon (iv) oxide "=0.112/22.4 \times 6.02 \times10^{23=3.01\times 10^21} molecules"